OpenStudy (anonymous):

What is the solution set of the following equation? -5x + x + 9 = -4x + 12

OpenStudy (anonymous):

-5x+x+9=-4x+12 group like terms -5x+x+4x=12-9 so if u solve them x=0 so answer is 0